OpenStudy (anonymous):

What is the sum of a 6–term geometric series if the first term is 20 and the last term is 336,140? 392,160 411,768 372,552 431,376

OpenStudy (anonymous):

\[\frac{6}{2}(20 + 336,140)\]

OpenStudy (anonymous):

1068.42

OpenStudy (saifoo.khan):

Hello Lahore! xD

OpenStudy (anonymous):

Lahori?

OpenStudy (anonymous):

saeed ajmal is from Lahore?

OpenStudy (saifoo.khan):

You are from Lahore. idk about Ajmal. im not not that much into cricket. Lahori was a typo. =)

OpenStudy (anonymous):

oh..no im from Karachi...

OpenStudy (saifoo.khan):

Wait. u said u are from Lahore! :O

OpenStudy (anonymous):

no i didnt...

OpenStudy (lilg132):

how did you get 1068.42 i dont get it

OpenStudy (saifoo.khan):

Lie detection!

OpenStudy (anonymous):

I have been there..but i am from karachi..

OpenStudy (saifoo.khan):

:O

OpenStudy (saifoo.khan):

Girls change their mind so often.. sign..

OpenStudy (anonymous):

Your prolly thinking of someone else

OpenStudy (lilg132):

i am from london now someone tell me please =(

OpenStudy (anonymous):

uh oh i read that mathematics series, sorry

OpenStudy (saifoo.khan):

sigh*

OpenStudy (saifoo.khan):

Nope, it was you!! i remember very clearly..

OpenStudy (anonymous):

why would I lie about being from lahore? pstch Karachi is like awesome..

OpenStudy (anonymous):

ugh auto-correct, arithmetic not mathematics

OpenStudy (anonymous):

geometric series

OpenStudy (anonymous):

Although, my grandfather was from India.I am half indian and half Paki...but more on Paki side cuz both my parents are from Karachi

OpenStudy (anonymous):

\[20 \times \frac{1 - r^7}{1 - r}\] \[r^5 = \frac{336,140}{20}\]

OpenStudy (anonymous):

16,807

OpenStudy (saifoo.khan):

Same here, my grandies are from India as well. =D Correct, Khi is Awesome!

OpenStudy (lilg132):

and its the opposite for me mine were from pakistan lol

OpenStudy (anonymous):

thts cool..lol @lilg132

OpenStudy (saifoo.khan):

LOL @lil

OpenStudy (anonymous):

Sorry Sum should be \[20 \times \frac{1 - r^6}{1 - r}\] Cool!@Saifoo, whatevs and lilg

OpenStudy (anonymous):

ugh I am confused please ignore me and my posts lilg and saifoo please help whatevs //something is wrong with me today :-/

OpenStudy (anonymous):

I am confused too..its ok Ishaan..thanks for helping...I'll just go with the first option..

OpenStudy (lilg132):

its 4 in the morning for me and i think that maths is getting to me ahaha see you all next time :p gdnightt

OpenStudy (anonymous):

goodnight..